calculate the wavelength and energy in joules for a single photon associated with 300 MHz signal of our new Oxford NMR. also, calculate the energy in kJ/mol for a mole of photons of this frequency.
Answer #1

energy of 1 mol = 6.022*10^23 * energy of 1 photon
= 6.022*10^23 * h*f
= 6.022*10^23 * (6.626*10^-34)*(300*10^6)
= 0.1197 J

Answer: 1.197*10^-4 KJ/mol